OpenStudy (anonymous):

i need help integer exponents 8-3^0=2^-1

OpenStudy (anonymous):

That statement doesnt work.. 3^0 is 1, anything to the exponent 0 is one. 2^-1 is just 1/2. so, its impossible that 8-1 could equal 1/2.. im not sure what your question is

OpenStudy (anonymous):

i meant + not =

OpenStudy (anonymous):

oh, okay, well, from what i said up there, you know that 3^0 is 1 (do you want an explanation of why?) and that anything to the power of -1 is just 1 over the number.. so, 8-1+1/2 :) i hope you can do the math from there :)

OpenStudy (anonymous):

no its ok i know that part and ok let me do it and tell you it

OpenStudy (anonymous):

do i do 8-1=7 so 7+1/2 so 7+2+9

OpenStudy (anonymous):

i mean 7+2=9

OpenStudy (anonymous):

I'm not certain about your last line... when adding fractions, so when adding 7+1/2, try to make both of the fractions have the same denominator. so, for example, 7 is the same as saying 14/2. so 14/2+1/2=15/2. remember that 7+1/2 isnt the same as 7+2 :)

OpenStudy (anonymous):

how do i add 7 onto a fraction 1/2

OpenStudy (anonymous):

 in step 1, i made 7 into a fraction over two, and then step two, because the two fractions have the same denominator, i just added the numerators